OpenStudy (anonymous):

Solve 4x^2+8x–252=0 by graphing

OpenStudy (tkhunny):

You should draw the graph. x = 0 corresponds to y = -252 Do more!

OpenStudy (anonymous):

4x^2+8x–252=0-> 0=4(x+1)^2-256 Vertex- x+1=0, x=-1 Axis of symmetry- -1 Domain- All real numbers Range- y>=4 Right?

OpenStudy (anonymous):

Oops, forgot to finish the vertex... Vertex- x+1=0, x=-1 y=4(-1+1)^2-256 y=4(0)^2-256 y=4(0)-256 y=-256 (-1,-256)

OpenStudy (tkhunny):

That's pretty good. There are a couple of refinements. 1) The Axis of Symmetry should be expressed as an equation of a line. x = -1 2) The Range is no good. For a parabola, this would always be a reference to the vertex. y >= -256

OpenStudy (anonymous):

Thank you (: